A liquid soap film in shape of a plane loop has an initial area 0.05 m2. If its area is slowly doubled then
the increase in its surface potential energy from its initial value will be (Surface tension of liquid = 0.2
N/m).

Grade:12th Pass

2 Answers

vikas askiitian expert
13 years ago

dE  = Tds                 

T = surface tension , ds change in area

     =(0.05)2(0.2) = 2*10-3 joules

SATYAM JAISWAL
5 years ago
U= Ts×Area
=2×0.2×0.05= 2×10-2J
 
This is correct answer . In this question we take a change in area and we multiplied by 2 because soap drop has two layer and surface tension is given so easily you will get answer.
 
Regards, 
Satyam Jaiswal